flor of a square room of side 10metres is to be completely covered with square tiles,each having length 50cm .the smallest number of tiles need is?
1
Expert's answer
2013-01-23T10:09:09-0500
Numberof tiles per 1 side:
n = 10 metres/50cm = 10/0.5 = 20
Therefore total numberof tiles:
N = n*n = 20*20 = 400
Answer: 400